OpenStudy (anonymous):

Jeannie has some $10 bills and some $20 bills.if she has 273 bills worth a total of $4370,how many of the bills are $10 bills and how many are $20 bills.?

hero (hero):

\(x=\) Number of ten dollar bills \(y = \) Number of twenty dollar bills \(x +y = 273\) \(10x + 20y = 4370\)

OpenStudy (anonymous):

Can you solved it for me

hero (hero):

Have you tried solving it?

OpenStudy (anonymous):

Yes But I am not familiar with this kind of equation thats why I need someone to help me

hero (hero):

Which are you more familiar with... Solving systems of equations by elimination or substitution?

OpenStudy (anonymous):

None am new in math

hero (hero):

To solve by elimination: Multiply both sides of the equation x+y=273 by 10: Then you'll have the following system: 10x+10y=2730 10x+20y=4370 Swap the rows to make it convenient for subtraction: 10x+20y=4370 10x+10y=2730 Subtract the second equation from the first to eliminate x variable: 10y=1640 Divide both sides by 10 y=164 Plug the value for y back in to the original equation to solve for x: The original equation was: x+y=273 If y=164 then x+164=273 x=273−164 x=109 So there are 109 ten dollar bills and 164 twenty dollar bills. If we check this we get: 109(10)+164(20)=4370